OpenStudy (anonymous):

a cube numbered 1-6 is tossed once, what is the probability that a number less than 4 shows on the top face of the cube?

OpenStudy (anonymous):

1/3 right?

OpenStudy (raden):

how many events of the number less than 4 ?

OpenStudy (anonymous):

3?

OpenStudy (anonymous):

and there is only one chance of it landing on the top face of the cube that why iguessed 1/3

OpenStudy (anonymous):

**that is why i guessed 1/3

OpenStudy (raden):

yeah, there are 3 events.we knowed that the space set of a cube = {1,2,3,4,5,6}, so the number all events = 6 now, use the formula of the probability of a event is : P(A) = numbers of events/numbers all events = 3/6 = 1/2

OpenStudy (anonymous):

oh i get it now, thank you..

OpenStudy (raden):

you're welcome